# rEGGression: an Interactive and Agnostic Tool for the Exploration of Symbolic Regression Models

### v1

@@ Abstract.
Regression analysis is used for prediction and to understand the effect of independent variables on dependent variables. Symbolic regression (SR) automates the search for non-linear regression models, delivering a set of hypotheses that balances accuracy with the possibility to understand the phenomena. Many SR implementations return a Pareto front allowing the choice of the best trade-off. However, this hides alternatives that are close to non-domination, limiting these choices. Equality graphs (e-graphs) allow to represent large sets of expressions compactly by efficiently handling duplicated parts occurring in multiple expressions. E-graphs allow to store and query all SR solution candidates visited in one or multiple GP runs efficiently and open the possibility to analyse much larger sets of SR solution candidates. We introduce rEGGression, a tool using e-graphs to enable the exploration of a large set of symbolic expressions which provides querying, filtering, and pattern matching features creating an interactive experience to gain insights about SR models. The main highlight is its focus in the exploration of the building blocks found during the search that can help the experts to find insights about the studied phenomena.This is possible by exploiting the pattern matching capability of the e-graph data structure.
@@
journalyear: 2024copyright: acmlicensedccs: Computing methodologies Symbolic and algebraic algorithmsccs: Mathematics of computing Genetic programming
@@ 1. Introduction
Regression analysis (Gelman, 2007) is an essential technique of machine learning and statistics with the objective of predicting unobserved data or, more importantly, gaining a better understanding of a given phenomena. In many situations, a linear relationship is assumed to give a rough estimation of the effects of the independent variables on the dependent variable. A linear model is easy to understand, as the effect is constant, but it may fail to capture the non-linear relationship of the variables. Using a non-linear model often requires expert knowledge about the phenomena and a manual and tedious process of exploring different recurrent patterns such as polynomials, rational functions, exponential growth, among others. Another possibility is to use a generic model such as generalized additive models (Hastie, 2017), which still can be considered interpretable or using opaque models such as neural networks (Bishop, 1994) and gradient boosting (Chen and Guestrin, 2016), which limits the analysis of the relationship of the variables.
@@
Symbolic regression (Koza, 1992; Kronberger et al., 2024a) automates this process while searching for a mathematical model of a system that approximates a collection of data points of interest. This technique can provide a valuable help for the task of scientific discovery (Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020).
The main highlight of symbolic regression is the possibility of finding a non-linear model accurately describing the data that still can be used for inference, data exploration, and interpretability, unlike opaque models.
@@
Although the main objective, when searching for a model, is to maximize accuracy, this comes with the cost of getting a complex model as a result. To alleviate this, many algorithms implement complexity control mechanisms such as penalization, model selection, maximum complexity constraint, and multi-objective optimization.
@@
Multi-objective optimization (Deb et al., 2002; Smits and Kotanchek, 2005; Kommenda et al., 2015) is particularly interesting because it transfers the decision on the best trade-off between accuracy and complexity to the user by presenting the whole Pareto front. Another possible variation is to return multiple solutions that behave similarly in the training set but differ when extrapolating (De Franca, 2023).
But in some situations the decision should be guided by some specificity of the system, that needs to be based on prior-knowledge of an expert (Haider and Kronberger, 2022; Martinek et al., 2024; Poursanidis et al., 2024).
@@
This prior-knowledge may come in the form of properties (Kronberger et al., 2022), unity obeying equations (Reuter et al., 2024), some desired building blocks that must be enforced (de França, 2022), or building blocks that should not appear in the expression (see for example, the constraints that can be created in (Cranmer, 2023)). For example, by inspecting the data, we may notice an oscillatory behavior similar to a damped oscillator, and then it is desired that the model follow the pattern $f(x;\theta)=g(x;\theta)e^{h(x;\theta)}\cos{(k(x;\theta))}$, where $g,h,k$ represents any function. Or, by further exploring the Pareto front, we observe a bias towards a recurrent pattern such as $\tanh{(g(x))}$ and we want to explore alternatives that do not contain it, even if having to use dominated expressions not appearing in the front. But, integrating prior-knowledge inside the search may not be trivial or can make the search space even harder to navigate. If we trust that the desired properties corresponds to an accurate model, we can explore the history of the visited expressions to find a better alternative. This, again, creates a burden to users as it is not trivial to explore $100$s of different expressions. Another insightful analysis is the frequent building blocks (Rosca, 1995; O’Reilly and Oppacher, 1995; Daida et al., 1999; McPhee et al., 2008) observed during the search that may reveal an important relationship or feature transformation that can be further inspected.
@@
In this paper, we propose rEGGression, an interactive tool that can help SR users to explore alternative models generated from different sources. These sources can be: the final population of a single run, the Pareto front, the entire history of visited expressions during the search, or a combination of those sources from multiple runs of the same or different algorithms. This can provide a rich library of alternative expressions generated from different biases, induced by different hyper-parameters or algorithms, that can bring invaluable information about the data. This tool supports simple queries such as querying for the top-N models filtered by size, complexity, and number of numerical parameters; insert and evaluate new expressions; list and evaluate sub-expressions of the already visited expressions, and also, more advanced queries such as calculate the frequency of common patterns (i.e., building blocks) observed in the set of models; and filter the expressions by patterns with a natural syntax.
@@
The main contribution of this paper is the pattern matching algorithm capable of efficiently searching and retrieving symbolic models that contain or not contain a certain pattern of interest. This is possible with the use of e-graph data structure (Willsey et al., 2021) created for the equality saturation algorithm, a technique used to alleviate the phase ordering problem in the optimization of computer programs during the compilation process. This technique was previously used in the context of symbolic regression in (de Franca and Kronberger, 2023; Kronberger and de França, 2024) to investigate the problem of unwanted overparameterization that increases the chance of a suboptimal fitting of the numerical parameters. The generated e-graph has another interesting feature that can be exploited by symbolic regression algorithms: it contains a database of patterns and equivalent expressions that can be easily matched against a new candidate expression.
@@
This paper is organized as follows. In Section 2 we will summarize the related works in exploration tools for symbolic regression. Section 3 will explain the basic concepts of equality saturation and the e-graph data structure. In Section 4 we will present rEGGression, showing the available commands, and giving additional details of its inner working. Finally, Section 5 gives some final remarks and expectations for the future.
@@ 2. Related work
Whenever we are confronted with multiple viable regression models with similar accuracy, we must eventually choose and commit with one of these alternatives to move forward with the regression analysis. This choice is often reduced to the accuracy-simplicity trade-off, in which we seek an accurate model that can still be interpreted up to a certain level. Model selection (Stoica and Selen, 2004) is part of the machine learning pipeline that deals with such a choice. Some model selection techniques that are often used are: evaluation of the model on a validation data to estimate the accuracy on unseen examples, criteria that takes simplicity into consideration such as BIC and AIC, the selection of the knee point of the Pareto front, or the recently proposed minimization of description length (Bartlett et al., 2023). While these criteria help to automate the process of model selection, they cannot accommodate additional constraints based on the expert knowledge. This prompts for an interactive approach that enables an efficient exploration of the alternative models.
@@
PySR (Cranmer, 2023) and PyOperon(Burlacu et al., 2020) both produce the Pareto front of the non-dominated expressions from the final GP generation.
Equality saturation has been used in the context of symbolic regression as a support tool to study the behavior of the search. de Franca and Kronberger  (de Franca and Kronberger, 2023; Kronberger and de França, 2024) show that many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters. This redundancy can increase the chance of failing to correctly optimize such parameters, leading to sub-optimal solutions. Simplification based on equality saturation was used to detect the equivalent expressions visited during the GP search (Kronberger et al., 2024b). As we will explain in the next section, the e-graph data structure and supporting structures used during equality saturation can be turned into an easy to use API providing ways to search through the model space with advanced queries and pattern matching of building blocks. This can simplify the exploration of alternative expressions while following an SQL-like domain specific language.
@@ 3. Equality saturation and e-graphs
These steps are illustrated in Fig. 2, where in Fig. 2a we have the initial expression, in which we match the rules $\alpha+\alpha\equiv 2\alpha$ and $\frac{\alpha}{\beta}\gamma\equiv\alpha\frac{\gamma}{\beta}$. These rules generates two new e-classes (Figs. 2b and 2c) that are merged afterwards (Fig. 2d). Although the idea is simple, it requires techniques from information retrieval for an efficient implementation of the matching procedure and maintenance of the structure (Willsey et al., 2021). One important characteristic of this structure and the algorithm is that it will always insert new information while keeping all previous states in the same structure.
@@
The e-class structure contains a set of e-nodes, a set of parent e-class ids with the corresponding e-nodes that points to the current e-class, and an extra field that contains information about the height of the e-class, the cost of the least complex expression that can be generated by that e-class, whether it evaluates to a constant value, and any additional information we may need for caching (e.g., positiveness, monotonicity, convexity, and measurement units information).
Finally, the database stores the information of recently created e-classes that must be analyzed and, eventually, merged, and a structure that stores the information of e-class ids for the smallest possible building blocks (e.g., a function or operator applied to anything). This database is a map of a token of our grammar (terminal or non-terminal) to a trie structure (Crochemore and Lecroq, 2009). This trie contains the information of all e-classes and their children containing that particular token. This allows us to efficiently retrieve all e-classes that have at least one e-node with a certain token and their corresponding children.
@@
Considering the e-class ids of Fig. 3 (small numbered boxes), querying the token $\times$ from the database will return a trie with the keys $4,6$, corresponding to the e-classes containing that token. The key $4$ will point to a trie with a single key $1$ which points to the key $2$. The key $6$ will point to the keys $1$ and $3$ corresponding to each e-node’s first child, and they will respectively point to the keys $4$ and $5$.
As an illustrative example, consider the query for the pattern $\alpha\times(\beta+\gamma)$. In the first step, it will create an empty list of substitution maps of tokens and variables to e-class ids:
@@
Next, it will retrieve the e-class ids for the token $\times$, which are $4,6$, creating two maps, one for each key:
@@
Following up the trie for the ids of the first childs, we replace the values for $\alpha$:
@@
Since $\alpha$ is a pattern variable, we do not need to check if the mapping is correct. The next entry for each trie will fill up the values for the $+$ token:
@@
Since both children are variables, the final map will be:
@@
Notice that if our pattern was $\alpha*(\beta+\beta)$, in the final step we would have
{×$\{\times\to 6,\alpha\to 3,+\to 4,\beta\to 2\}$
after checking whether that both children of $+$ points to $2$.
@@
After the final step, this procedure will return a set of substitution maps that will point to each match of that pattern. In our example, if we want to create the equivalent expression $(\alpha\times\beta)+(\alpha\times\gamma)$, we start by inserting $(\alpha\times\beta)$, or the e-node (×$(\times,[3,2])$ (Fig.4a), which would generate a new e-class with id $7$, we insert the e-node corresponding to the right side, (×$(\times,[3,2])$, which would return the now existent e-class id $7$, and we would finally create the e-node $(+,[7,7])$, which would prompt the creation of a new e-class id $8$ (Fig.4b). This e-class would be marked to be merged with the e-class id $6$ as they are equivalent.
Regarding the complexity of these operations, the building of the database is linear, while the querying is bounded by $O(\sqrt{qn^{m}})$, where $q$ is the number of returned queries, $n$ is the number of e-nodes in the e-graph, and $m$ is the number of conjunctive queries, proportional to the number of non-terminal nodes.
More details about an efficient implementation of e-graph and equality saturation can be found in (Willsey et al., 2021) and for the pattern matching algorithm in (ZHANG et al., 2021).
There are currently a few implementations of equality saturation, notably egg [^4] https://docs.rs/egg/latest/egg/ egg stands for e-graphs good in Rust (Willsey et al., 2021), hegg [^5] https://github.com/alt-romes/hegg/ in Haskell, Metatheory.jl [^6] https://juliasymbolics.github.io/Metatheory.jl/dev/egraphs/ in Julia.
@@
In this paper, we are proposing a new software, called rEGGression, an exploration and support tool for symbolic regression models that uses e-graphs to perform advanced queries. This tool is compatible with the output of many popular symbolic regression implementations.
The key highlights of rEGGression are:
@@
The implementation of rEGGression follows closely the implementation of e-graph and equality saturation described in (Willsey et al., 2021) and the pattern matching in (ZHANG et al., 2021). The name is a reference to the current main e-graph library egg written in Rust [^7] https://egraphs-good.github.io/.

### v2

@@ Abstract.
Regression analysis is used for prediction and to understand the effect of independent variables on dependent variables. Symbolic regression (SR) automates the search for non-linear regression models, delivering a set of hypotheses that balances accuracy with the possibility to understand the phenomena. Many SR implementations return a Pareto front allowing the choice of the best trade-off. However, this hides alternatives that are close to non-domination, limiting these choices. Equality graphs (e-graphs) allow to represent large sets of expressions compactly by efficiently handling duplicated parts occurring in multiple expressions. The e-graphs allow to efficiently store and query all solution candidates visited in one or multiple runs of different algorithms and open the possibility to analyze much larger sets of SR solution candidates. We introduce rEGGression, a tool using e-graphs to enable the exploration of a large set of symbolic expressions which provides querying, filtering, and pattern matching features creating an interactive experience to gain insights about SR models. The main highlight is its focus in the exploration of the building blocks found during the search that can help the experts to find insights about the studied phenomena. This is possible by exploiting the pattern matching capability of the e-graph data structure.
@@
journalyear: 2025copyright: acmlicensedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25), July 14–18, 2025, Malaga, Spaindoi: 10.1145/3712256.3726385isbn: 979-8-4007-1465-8/2025/07ccs: Computing methodologies Symbolic and algebraic algorithmsccs: Mathematics of computing Genetic programming
@@ 1. Introduction
Regression analysis (Gelman, 2007) is an essential technique of machine learning and statistics with the objective of predicting unobserved data or, more importantly, gaining a better understanding of a given phenomena. In many situations, a linear relationship is assumed to give a rough estimation of the effects of the independent variables on the dependent variable. A linear model is easy to understand, as the effect is constant, but it may fail to capture the non-linear relationship of the variables. Using a non-linear model often requires expert knowledge about the phenomena and a manual and tedious process of exploring different recurrent patterns such as polynomials, exponential growth, among others. Another possibility is to use a generic model such as generalized additive models (Hastie, 2017), which still can be considered interpretable or using opaque models such as neural networks (Bishop, 1994) and gradient boosting (Chen and Guestrin, 2016), which limits the analysis of the relationship of the variables.
@@
Symbolic regression (Koza, 1992; Kronberger et al., 2024a) (SR) automates this process by searching for a mathematical model of a system that approximates a collection of data points of interest. This technique can provide a valuable help for the task of scientific discovery (Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020). The main highlight of symbolic regression is the possibility of finding a non-linear model accurately describing the data that still can be used for inference, data exploration, and interpretability.
@@
Although the main objective, when searching for a model, is to maximize accuracy, this comes with the cost of getting a complex model as a result. For this reason, many algorithms implement complexity control mechanisms such as penalization, model selection, maximum complexity constraint, and multi-objective optimization.
@@
Multi-objective optimization (Deb et al., 2002; Smits and Kotanchek, 2005; Kommenda et al., 2015) transfers the decision on the best trade-off between accuracy and complexity to the user by presenting the whole Pareto front. Another alternative is to return multiple solutions that are semantically similar int the training set, but different outside that boundaries (De Franca, 2023). But in some situations the decision should be guided by some specificity of the system, that needs to be based on prior-knowledge of an expert (Haider and Kronberger, 2022; Martinek et al., 2024; Poursanidis et al., 2024).
@@
This prior-knowledge may be properties (Kronberger et al., 2022), unity obeying equations (Reuter et al., 2024), desired building blocks that must be enforced (de França, 2022) or that should not appear in the expression (see for example, the constraints that can be created in (Cranmer, 2023)). For example, by inspecting the data, we may notice an oscillatory behavior similar to a damped oscillator, and then it is desired that the model follow the pattern $f(x;\theta)=g(x;\theta)e^{h(x;\theta)}\cos{(k(x;\theta))}$, where $g,h,k$ represents any function. Or, by further exploring the Pareto front, we observe a bias towards a recurrent pattern such as $\tanh{(g(x))}$ and we want to explore alternatives that do not contain it, even if having to use dominated expressions not appearing in the front. But, integrating prior-knowledge inside the search may not be trivial or can make the search space even harder to navigate. If we trust that the desired properties corresponds to an accurate model, we can explore the history of the visited expressions to find an accurate model that behaves as expected. This, again, creates a burden to users as it is not trivial to explore hundreds of different expressions. Another insightful analysis is the frequent building blocks (Rosca, 1995; O’Reilly and Oppacher, 1995; Daida et al., 1999; McPhee et al., 2008) observed during the search that may reveal an important relationship or feature transformation that can be further inspected.
@@
In this paper, we propose rEGGression, an interactive tool that can help SR users to explore alternative models generated from different sources. These sources can be: the final population of a single run, the Pareto front, the entire history of visited expressions during the search, or a combination of those sources from multiple runs of the same or different algorithms. This can provide a rich library of alternative expressions generated departing from different priors (i.e., hyper-parameters, algorithms, random seeds), that can bring invaluable information about the data. This tool supports simple queries such as querying for the top-N models filtered by size, complexity, and number of numerical parameters; insert and evaluate new expressions; list and evaluate sub-expressions of the already visited expressions, and also, more advanced queries such as calculate the frequency of common patterns (i.e., building blocks) observed in the set of models; and filter the expressions by patterns with a natural syntax.
@@
The unique features of this tool is the pattern matching capabilities and the analysis of building blocks. The pattern matching allows to filter the current set of expressions to those that contains (or not contain) a certain functional pattern, this allows to explore models based on domain knowledge or additional assumptions that could not be incorporated during the search. The analysis of the building blocks reports the most frequent building blocks explored during the search or the building blocks with the largest average fitness. This information brings additional insights of possible patterns that contributes to obtaining a high accurate model, providing new insights about the observations. The interactive nature of this tool lies in the ability to refine the search and shifting the focus in different regions of the space of available expressions. These unique features are possible with the use of e-graph data structure (Willsey et al., 2021) created for the equality saturation algorithm, a technique used to alleviate the phase ordering problem in the optimization of computer programs during the compilation process. This technique was previously used in the context of symbolic regression in (de Franca and Kronberger, 2023; Kronberger and de França, 2024) to investigate the problem of unwanted overparameterization that increases the chance of a suboptimal fitting of the numerical parameters. The generated e-graph has another interesting feature that can be exploited by symbolic regression algorithms: it contains a database of patterns and equivalent expressions that can be easily matched against a new candidate expression.
@@
This paper is organized as follows. In Section 2 we will summarize the related works in exploration tools for symbolic regression. Section 3 will explain the basic concepts of equality saturation and the e-graph data structure. In Section 4 we will present rEGGression, showing the available commands, and giving additional details of its inner working, specially how the pattern matching and building blocks exploration algorithms work. Finally, Section 5 gives some final remarks and expectations for the future.
@@ 2. Related work
Whenever we are confronted with multiple viable regression models with similar accuracy, we must eventually choose and commit with one of these alternatives to move forward with the analysis. This choice is often reduced to the accuracy-simplicity trade-off, in which we seek an accurate model that can still be interpreted up to a certain level. Model selection (Stoica and Selen, 2004) is part of the machine learning pipeline that deals with such a choice. Some common model selection techniques comprehends the evaluation of the model on a validation data to estimate the accuracy on unseen examples, criteria that takes simplicity into consideration such as BIC and AIC, the selection of the knee point of the Pareto front, or the recently proposed minimization of description length (Bartlett et al., 2023). While these criteria help to automate the process of model selection, they cannot accommodate additional constraints based on the domain knowledge. This prompts for an interactive approach that enables an efficient exploration of the many alternative models that might be available.
@@
Equality saturation has been used in the context of symbolic regression as a support tool to study the behavior of the search. de Franca and Kronberger  (de Franca and Kronberger, 2023; Kronberger and de França, 2024) show that many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters that can lead to sub-optimal results. Simplification based on equality saturation was used to detect the equivalent expressions visited during the GP search (Kronberger et al., 2024b). As we will explain in the next section, the e-graph data structure and supporting structures used during equality saturation can be turned into an easy to use API providing ways to search through the model space with advanced queries and pattern matching of building blocks. This can simplify the exploration of alternative expressions while following an SQL-like domain specific language.
@@ 3. Equality saturation and e-graphs
Steps $1$ to $3$ for a single equivalence rule are illustrated in Fig. 2, where in Fig. 2a we have the initial expression, in which we match the rules $\alpha+\alpha\equiv 2\alpha$ and $\frac{\alpha}{\beta}\gamma\equiv\alpha\frac{\gamma}{\beta}$. These rules generates two new e-classes (Figs. 2b and 2c) that are merged afterwards (Fig. 2d). Steps $1$ and $2$ are repeated for every equivalence rule and matched pattern and step $3$ is performed only after all new e-classes are recreated from the previous steps. Reaching saturation is often impossible as some rules can lead to an infinite sized e-graph. Although the idea is simple, it requires techniques from information retrieval for an efficient implementation of the matching procedure and maintenance of the structure (Willsey et al., 2021). One important characteristic of this structure and the algorithm is that it will always insert new information while keeping all previous states in the same structure.
@@
The e-class structure contains a set of e-nodes, a set of parent e-class ids with the corresponding e-nodes that points to the current e-class, and an extra field that contains information about the height of the e-class, the cost of the least complex expression that can be generated by that e-class, whether it evaluates to a constant value, and any additional information we may want to store (e.g., positiveness, monotonicity, convexity, and measurement units information).
Finally, the database stores the information of recently created e-classes that must be analyzed and merged, and a structure that stores the information of e-class ids for the smallest possible building blocks (e.g., a function or operator applied to anything). This database is a map of a token of our grammar (terminal or non-terminal) to a trie structure (Crochemore and Lecroq, 2009). This trie contains the information of all e-classes and their children containing that particular token. This allows us to retrieve all e-classes that have at least one e-node with a certain token and their corresponding children. Considering the e-class ids of Fig. 3 (small numbered boxes), querying the token $\times$ from the database will return a trie with the keys $4,6$, corresponding to the e-classes containing that token. The key $4$ will point to a trie with a single key $1$ which points to the key $2$. The key $6$ will point to the keys $1$ and $3$ corresponding to each e-node’s first child, and they will respectively point to the keys $4$ and $5$.
@@
As an illustrative example, consider the query for the pattern $\alpha\times(\beta+\gamma)$. In the first step, it creates an empty list of substitution maps of tokens and variables to e-class ids:
@@
Next, it retrieves the e-class ids for the token $\times$, which are $4,6$, creating two maps, one for each key:
@@
Following up the trie for the ids of the first child, it replaces the values of $\alpha$:
@@
Since $\alpha$ is a pattern variable, we do not need to check whether the mapping is correct. The next entry for each trie will fill up the values for the $+$ token:
@@
Since both children are pattern variables, the final map will be:
@@
After the final step, this procedure will return a set of substitution maps that will point to each match of that pattern and can be used to create equivalent expressions. Suppose we want to create the equivalent expression $(\alpha\times\beta)+(\alpha\times\gamma)$, we start by inserting $(\alpha\times\beta)$, or the e-node (×$(\times,[3,2])$ (Fig.4a), which would generate a new e-class with id $7$, we then insert the e-node corresponding to the right side of the expression, (×$(\times,[3,2])$, which would simply return the e-class id $7$ (i.e., it will not create the same expression again), and we would finally create the e-node $(+,[7,7])$, which would prompt the creation of a new e-class id $8$ (Fig.4b). This e-class would be marked to be merged with the e-class id $6$ as they are equivalent.
Regarding the complexity of these operations, the building of the database is linear, while the querying is bounded by $O(\sqrt{qn^{m}})$, where $q$ is the number of returned queries, $n$ is the number of e-nodes in the e-graph, and $m$ is the number of conjunctive queries, proportional to the number of non-terminal nodes. More details of an efficient implementation of e-graph and equality saturation can be found in (Willsey et al., 2021) and for the pattern matching algorithm in (ZHANG et al., 2021).
@@
In this paper, we are proposing a new software, called rEGGression, an exploration tool for symbolic regression models that uses e-graphs to perform advanced queries. This tool is compatible with the output of many popular symbolic regression implementations (see Sec.4.10). The key highlights of rEGGression are:
@@
The implementation of rEGGression follows closely the implementation of e-graph and equality saturation described in (Willsey et al., 2021) and the pattern matching in (ZHANG et al., 2021). The name of this tool is a direct reference to the current main e-graph library egg written in Rust [^4] https://egraphs-good.github.io/.